# AdaViPro: Region-based Adaptive Visual Prompt for Large-Scale Models Adapting

### v1

@@ 1 Introduction
Recently, fuelled by the label-free self-supervised learning, Transformer [1] and its variants have demonstrated remarkable success in a wide range of areas, such as natural language processing [2, 3], computer vision [4, 5, 6] and audio signal processing [7]. However, as the scale of the model size grows, the conventional pretraining-finetuning paradigm becomes impractical [8, 9]. Fully fine-tuning these large pre-trained models for varied downstream tasks proves to be expensive and inflexible. Meanwhile, due to the limited training data of downstream tasks, improper fine-tuning might compromise the universal representations established during pre-training, diminishing their effectiveness.
@@ 2.1 Large-scale Models
With the advancement of self-supervised learning [2, 15, 16] and the availability of large-scale datasets, the problem of ‘data hunger’ in high-capacity deep learning models has been effectively addressed. In this context, the scalability potential of the transformer [1] is continuously explored, leading to the development of large-scale models in NLP that exhibit formidable capabilities [3]. This success has also been mirrored in the field of computer vision [4, 16] and multi-modal learning [17, 18]. Recent works propose to pre-train large-scale foundational visual models via the self-supervised learning pipeline, including methods like contrastive learning [15, 17] and masked image modeling [16], among others. Our work is orthogonal to the pretraining of foundational models. We focus on how to transfer the generalized representations of large pre-trained models to downstream tasks.
@@ 3.1 Preliminaries
In this paper, we build AdaViPro based on the prompt method VP [10], which is a kind of model-agnostic universal pixel-level prompt. Note that the techniques introduced in this paper are versatile and can also be applied to other prompt methods. To provide legibility and context, we first review the main idea of VP. Specifically, taking CLIP as an example, let $f\left(\cdot\mid\theta_{v}\right)$ and $f\left(\cdot\mid\theta_{t}\right)$ denote the vision and text encoder respectively. Given a category collection $C=\{c_{1},c_{2},...,c_{K}\}$ and an image-text pair $M=\{x_{v},c_{v}\}$, VP can be formulated as:
@@ 4.2 Comparison with Baseline Methods
As indicated in Table 1, we first compare the performance of AdaViPro with several baselines across 9 datasets, utilizing the VP’s [10] default prompt size 30. The results demonstrate that AdaViPro outperforms TP and VP across all datasets by a considerable margin, with an average improvement of 2.2%. Specifically, in the CIFAR10 and EuroSAT, AdaViPro exceeded the performance of Linear fine-tuning, achieving impressive accuracy rates of 95.4% and 96.8%, respectively. By strategically incorporating ‘where to add’ into optimization, AdaViPro generates adaptive prompts that avoid obscuring objects within the image, thereby further enhancing recognition performance.
@@
With the prompt size of 112, the accuracy rates notably dropped to 38.4% and 54.5% in DTD and UCF101, respectively. Similar results can also be observed in Figure 2 as well. VP’s performance undergoes a steep decline with the increase in prompt size. Larger prompt sizes and an increased number of tunable parameters do not lead to performance improvements; rather, they result in the loss of key information in the input images. In contrast, AdaViPro consistently maintains robust performance, even achieving the best results in DTD with a prompt size of 112. This demonstrates that adaptive prompts enable instance-specific generation, effectively avoiding the obstruction of objects and information within the image. Number of Tunable Parameters. We compare the tunable parameters number of AdaViPro with different baseline methods, as shown in Table 3. Fully fine-tuning requires adjusting all parameters of the model, which is quite expensive. The number of tunable parameters for linear fine-tuning and VP is roughly the same, but VP still exhibits a significant performance gap compared to the former. Our AdaViPro adds only 0.13M additional parameters, enabling the generation of adaptive prompts, which achieves a well-balanced trade-off between accuracy and parameter efficiency.

### v2

@@ 1 Introduction
Recently, fuelled by the label-free self-supervised learning, Transformer [1] and its variants have demonstrated remarkable success in a wide range of areas, such as natural language processing [2, 3], computer vision [5, 6] and audio signal processing [7]. However, as the scale of the model size grows, the conventional pretraining-finetuning paradigm becomes impractical [8, 9]. Fully fine-tuning these large pre-trained models for varied downstream tasks proves to be expensive and inflexible. Meanwhile, due to the limited training data of downstream tasks, improper fine-tuning might compromise the universal representations established during pre-training, diminishing their effectiveness.
@@ 2.1 Large-scale Models
With the advancement of self-supervised learning [2, 16] and the availability of large-scale datasets, the problem of ‘data hunger’ in high-capacity deep learning models has been effectively addressed. In this context, the scalability potential of the transformer [1] is continuously explored, leading to the development of large-scale models in NLP that exhibit formidable capabilities [3]. This success has also been mirrored in the field of computer vision [16] and multi-modal learning [17, 18]. Recent works propose to pre-train large-scale foundational visual models via the self-supervised learning pipeline, including methods like contrastive learning [17] and masked image modeling [16], among others. Our work is orthogonal to the pretraining of foundational models. We focus on how to transfer the generalized representations of large pre-trained models to downstream tasks.
@@ 3.1 Preliminaries
In this paper, we build AdaViPro based on the prompt method VP [10], which is a kind of model-agnostic universal pixel-level prompt. Note that the techniques introduced in this paper are versatile and can also be applied to other prompt methods. To provide legibility and context, we first review the main idea of VP. Specifically, taking CLIP as an example, let $f\left(\cdot\mid\theta_{v}\right)$ and $f\left(\cdot\mid\theta_{t}\right)$ denote the vision and text encoder respectively. Given the category collection $C=\{c_{1},c_{2},...,c_{K}\}$ and image-text pair $M=\{x_{v},c_{v}\}$, VP can be formulated as:
@@ 4.2 Comparison with Baseline Methods
As indicated in Table 1, we first compare the performance of AdaViPro with several baselines across 9 datasets, utilizing the VP’s [10] default prompt size 30. The results demonstrate that AdaViPro outperforms TP and VP across all datasets by a considerable margin, with an average improvement of 2.2%. Specifically, in the CIFAR10 and EuroSAT, AdaViPro exceeded the performance of Linear fine-tuning, achieving impressive accuracy rates of 95.4% and 96.8%, respectively. By strategically incorporating ‘where to add’ into optimization, AdaViPro generates adaptive prompts that avoid obscuring objects, enhancing the recognition performance.
@@
With the prompt size of 112, the accuracy rates notably dropped to 38.4% and 54.5% in DTD and UCF101, respectively. Similar results can also be observed in Figure 2 as well. VP’s performance undergoes a steep decline with the increase in prompt size. Larger prompt sizes and an increased number of tunable parameters do not lead to performance improvements; rather, they result in the loss of key information in the input images. In contrast, AdaViPro consistently maintains robust performance, even achieving the best results in DTD with a prompt size of 112. This demonstrates that adaptive prompts enable instance-specific generation, effectively avoiding the obstruction of objects and information within the image.
Number of Tunable Parameters. We compare the tunable parameters number of AdaViPro with different baseline methods, as shown in Table 3. Fully fine-tuning requires adjusting all parameters of the model, which is quite expensive. The number of tunable parameters for linear fine-tuning and VP is roughly the same, but VP still exhibits a significant performance gap compared to the former. Our AdaViPro adds only 0.13M additional parameters, enabling the generation of adaptive prompts, which achieves a well-balanced trade-off between accuracy and parameter efficiency.